Quiz: Chapter 10 Quiz Time Remaining: 00:59:21 Submit Quiz This Question: 1 pt 5 of 5 (0 complete) This Quiz: 5 pts possible Construct a scatter plot and identify the mathematical model that best fils the given data. Assume that the model is to be used only for the scope of the given data, and consider only linear, quadratic, logarithmic, exponential, and power models. 4.9 7.8 10.7 13.6 16.5 19.4 What is the equation of the best model? Select the most appropriate choice below and fill in the answer box(es) to completa your choica. (Round to three decimal places as needed.) O D. In x Click to select and enter your answers).
Answer #1

data

x y
1 4.9
2 7.8
3 10.7
4 13.6
5 16.5
6 19.4

Scatter-plot

25 20 15 10 0 0 4

This perfect straight line

hence

y =a +bx is correct form

y = 2 + 2.9 x

option C) y = 2 + 2.9*x is correct